OpenStudy (anonymous):

How do you write the equation in standard form, then determine the vertex, axis and direction in which each parabola opens? y^2 + 2y - 4x - 3 = 0

OpenStudy (anonymous):

y=mx+b

myininaya (myininaya):

so we want to write this in x=a(y-h)^2+k form so solve for x giving us x=[y^2+2y-3]/4 x=y^2/4+y/2-3/4 x=1/4(y^2+2y)-3/4 x=1/4(y^2+2y+1)-3/4-1/4 x=1/4(y+1)^2-1

OpenStudy (anonymous):

Thank you. My pc is acting up, so I didn't even know you messaged.

OpenStudy (anonymous):

Thank you. My pc is acting up, so I didn't even know you messaged.